5+1/5+1 1/3=
7 1/5 4 1/3=
1/2x5 5/8 x4=
5/8 divided 3/4=
x/6=5/10
I REALLY NEED HELP!!!! DONT REMEMBER ANY OF THIS.PLZ HELP.

Answer by sudhanshu_kmr(1152)


5+1/5+1 1/3=
= 5 + 1/5 + 4/3
= [75+ 3 +20]/15
= 98/15



5/8 divided 3/4
= 5/8 * 4/3
= 5/6




x/6=5/10
=> x = 5/10 * 6
=> x = 1/2 *6
=> x = 3


if any doubt, contact..